# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v3

@@ 15         Record V-I near-miss, set $t_{c}$;
While this geometric framework enables precise localization of near misses, it does not capture the probabilistic structure or temporal evolution of COR. To address this, the extracted blockwise minima of 2D-TTC are modeled using the EVT model. For each block, the model estimates the probability that 2D-TTC falls below a critical safety threshold, which defines the COR metric as a short-term indicator of crash risk. Block duration is chosen to balance temporal sensitivity with statistical stability across diverse contexts. This framework, therefore, combines deterministic near-miss detection with probabilistic EVT modeling to enable localized, data-driven assessment of crash-prone conditions. The complete methodological pipeline, from trajectory simulation to dynamic risk estimation, is summarized in Fig. 2.5.
@@ 3 Extreme value theory (EVT)
Extreme Value Theory (EVT) provides the statistical foundation for estimating COR by extrapolating rare, high-severity events from frequent near-misses. In this study, extreme events are represented by block maxima of 2D-TTC derived from high-resolution vehicle trajectories (Algorithm 1). EVT allows risk estimation without requiring large volumes of historical crash data (Zheng and Sayed, 2020; Fu and Sayed, 2022a). Two EVT approaches are commonly applied: the Block Maxima (BM) method, which fits maxima or minima from fixed-duration blocks using the Generalized Extreme Value (GEV) distribution (Coles et al., 2001), and the Peaks Over Threshold (POT) method, which models threshold exceedances using the Generalized Pareto Distribution (GPD). Although POT can be data-efficient, it is highly sensitive to threshold choice and less compatible with discrete sensor-derived datasets. By contrast, BM aligns naturally with real-time AV data, avoids subjective thresholds, and supports scalable, time-series analysis (Fu and Sayed, 2022a; Songchitruksa and Tarko, 2006). Consequently, this study adopts the BM–GEV framework (see Fig. 2.5).
The GEV has been widely applied in traffic safety research. Early studies modeled near-missets with univariate and multivariate GEV formulations (Orsini et al., 2020; Zhang et al., 2019; Zheng et al., 2014). Recent advances incorporate nonstationary structures, linking GEV parameters to covariates such as vehicle dynamics (Anis et al., 2025b; Fu and Sayed, 2021c, 2022b). Bayesian hierarchical models have also addressed sparse data and unobserved heterogeneity across vehicles, sites, and traffic contexts. The increasing availability of high-resolution AV trajectory and HD map data has highlighted the need for flexible GEV formulations that jointly capture V–V and V–I interactions.
Despite this progress, most UGEV models still neglect the joint influence of roadway geometry and vehicle dynamics, two critical but distinct sources of heterogeneity. Roadway geometry (e.g., lane count, curvature, median type) is spatially structured and relatively stable within segments or intersections, while vehicle dynamics (e.g., speed, acceleration) vary temporally and spatially, driving short-term behavioral volatility. Capturing this multilevel heterogeneity is crucial for estimating corridor and network level crash risk. To address this gap, this study develops a hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) UGEV model. Unlike standard random-parameter models, which assign variation independently to each unit, grouped random parameters capture shared variation across clusters (e.g., segment types, corridor directions) while allowing for within-group variability. This structure provides both flexibility and parsimony, making it well-suited for high-resolution, corridor-scale crash risk modeling (Washington et al., 2020; Cai et al., 2018; Islam et al., 2023).
@@ 3.1 Univariate generalized extreme value (UGEV)
As established in the previous section, the BM-based UGEV model is employed to characterize the tail behavior of 2D-TTC, which represents the most critical near-miss events in both V–V and V–I interactions. To implement this framework, continuous vehicle trajectories are partitioned into fixed-duration windows (e.g., 11 s). Within each window, the minimum 2D-TTC is extracted and negated, consistent with the upper-tail modeling convention of UGEV (Coles et al., 2001). The resulting sequence of BM serves as the input for parameter estimation. Throughout this study, we focus on maxima, denoted as $M_{n}$, under the assumption that each block consists of $n$ independent and identically distributed (i.i.d.) random variables, $X_{1},X_{2},\ldots,X_{n}$. To align with conventional UGEV formulations, we consider the negative values of the original variables derived from Algorithm 1, thereby defining the block maximum as $M_{n}=\max\{-X_{1},-X_{2},\ldots,-X_{n}\}$. Under suitable regularity conditions, if there exist sequences of constants $a_{n}\in\mathbb{R}$ and $b_{n}>0$, the normalized block maxima $M_{n}^{*}=(M_{n}-a_{n})/b_{n}$ converge in distribution to a non-degenerate limit $G(x)$ as $n\to\infty$, then: $\Pr\left(\frac{M_{n}-a_{n}}{b_{n}}\leq X\right)\xrightarrow[n\to\infty]{}f(x)$, where $G(x)$ is a non-degenerate distribution function. According to EVT, the limiting distribution $G(x)$ must belong to one of three families: Gumbel, Fréchet, or Weibull, each of which can be captured by the parametric form of the GEV distribution (Coles et al., 2001). Let the UGEV be defined as:
@@
3.1.1 Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model
To capture structured heterogeneity in extreme outcomes across the corridor, we employ a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model. Unlike conventional random-parameter models that assign variation to individual units, the HBSGRP formulation introduces variation at the group level. This structure allows systematic differences across predefined observational groups while maintaining parsimony and interpretability. In this specification, covariates are divided into two sets: fixed effects $X$, which apply uniformly across all observations, and group-varying effects $Z$, whose coefficients differ across groups. For a dataset divided into $K$ groups, with observations $i$ belonging to group $k(i)$, both the intercept and selected slopes are allowed to vary by group. This enables the capture of localized relationships while preserving overall model coherence. The model is expressed within a three-layer hierarchical Bayesian framework. The data layer specifies the likelihood of observed block maxima of 2D-TTC from V–V and V–I interactions, assuming a UGEV distribution. The process layer links covariates to UGEV parameters through fixed and group-level random effects, with optional latent terms (e.g., Gaussian processes) to capture residual spatial–temporal dependence. The prior layer defines prior distributions for all parameters and hyperparameters, supporting regularization and uncertainty quantification.
@@ 3.1.1 Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model
In this formulation, $q(\Theta\mid X)$ denotes the posterior distribution, integrating both the data likelihood and prior knowledge. The term $q_{\text{data}}(X\mid\Theta)$ corresponds to the likelihood function defined by the EVT assumptions. The component $q_{\text{process}}(\Theta\mid\Psi)$ captures the hierarchical dependency structure, including the group-varying parameters. Lastly, $q_{\text{prior}}(\Psi)$ encompasses the prior distributions of hyperparameters that control the model’s flexibility and penalize over-complexity.
@@
$q_{\text{data}}(X_{k,i}\mid\Theta)=\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}\frac{1}{\sigma_{k,i}}\exp\left\{-\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1-1/\xi_{k,i}}$ (23)
The location ($\mu_{k,i}$), scale ($\sigma_{k,i}$), and shape ($\xi_{k,i}$) parameters are expressed as functions of group-level fixed-effect covariates $X_{k,j}$ (e.g., road geometry, speed limit) and sample-specific random-effect covariates $W_{k,i,j}$ (e.g., speed, acceleration, time-to-collision) as follows:
$\mu_{k,i}=\beta_{\mu,0}+\sum_{j=1}^{J}\beta_{\mu,j}X_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\mu,j,k}W_{k,i,j}$ (24)
$\vartheta_{k,i}=\beta_{\sigma,0}+\sum_{j=1}^{J}\beta_{\sigma,j}X_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\sigma,j,k}W_{k,i,j},\quad\sigma_{k,i}=\exp(\vartheta_{k,i})$ (25)
$\xi_{k,i}=\beta_{\xi,0}+\sum_{j=1}^{J}\beta_{\xi,j}X_{k,j}$ (26)
In the above equations, $X_{k,j}$ defined as a group-level fixed-effect covariate $j$, $W_{k,i,j}$is a sample-specific random-effect covariate $j$. $\beta_{\mu,0},\beta_{\sigma,0},\beta_{\xi,0}$ as fixed intercepts for respective parameters. $\beta_{\mu,j},\beta_{\sigma,j},\beta_{\xi,j}$: fixed-effect coefficients. $\gamma_{\mu,j,k},\gamma_{\sigma,j,k}$: random-effect coefficients, varying across groups $k$. Therefore, the likelihood for the process layer, modeling within-group variability, is defined as:
@@
where $\bar{\mu}_{k}$, $\bar{\vartheta}_{k}$, and $\bar{\xi}_{k}$ are group-specific average parameters.
@@
where $\text{IG}(\alpha,\beta)$ denotes an inverse gamma prior distribution ensuring positivity constraints for variance parameters.
@@ 3.2 COR estimation
Building on the HBSGRP–UGEV estimates, the framework computes COR by evaluating the probability that 2D-TTC for both V–V and V–I interactions falls below a critical threshold within each block. Unlike severity-stratified approaches, this framework estimates the total number of latent extreme near-miss events regardless of outcome severity, allowing for a holistic assessment of potential COR under mixed traffic conditions.
Let $k,i$ denote the block $i$ within the group $k$, where groups correspond to spatial or functional corridor classifications (e.g., intersections or directional segments). The minimum 2D-TTC in each block is modeled with the HBSGRP-UGEV distribution, parameterized by $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$. Let $\omega$ be the critical 2D-TTC threshold below which a near-miss is considered crash-prone. The probability of a crash-level event in block $i$ of group $k$ is:
$P_{k,i}^{\text{crash}}=\Pr(2D-TTC_{k,i}\leq\omega)=1-\exp\left\{-\left[1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\},\quad\text{where }1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)>0$ (30)
Assuming a homogeneous Poisson process for exceedances, the expected COR for the group $k$ is obtained by aggregating exceedance probabilities across blocks and normalizing by exposure duration $T$. Here $y_{k,i}$ is the number of near-misses (V-V or V-I) in block $i$, and $N_{k}$ the number of blocks in group $k$. The expected COR for group $k$ is:
$COR_{k}=\frac{1}{T}\sum_{i=1}^{N_{k}}y_{k,i}\cdot P_{k,i}^{\text{crash}}=\frac{1}{T}\sum_{i=1}^{N_{k}}y_{k,i}\cdot\left[1-\exp\left\{-\left[1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\right]$ (31)
The total expected crash frequency across the study area is then:
$CF_{\text{total}}=\sum_{k=1}^{K}COR_{k}=\frac{1}{T}\sum_{k=1}^{K}\sum_{i=1}^{N_{k}}y_{k,i}\cdot\left[1-\exp\left\{-\left[1+\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\right]$ (32)
This formulation integrates exceedance probabilities from the HBSGRP–UGEV model with observed V–V and V–I near-miss counts, maintaining a direct link between real-world interaction intensity and modeled crash propensity. By separating probability from exposure, the framework yields interpretable, proactive crash frequency estimates aligned with EVT-based safety assessment.

### v4

@@ 15         Record V-I near-miss, set $t_{c}$;
While the above procedure deterministically detects and localizes near-misses, COR requires a probabilistic mapping from near-miss severity to short-term crash risk. Accordingly, BM, defined as the minimum 2D-TTC observed across the selected window, is modelled using EVT. COR is then defined as the estimated probability that 2D-TTC falls below a pre-specified critical threshold. Here, this block definition preserves short-lived maneuver dynamics at the native trajectory resolution while yielding a sufficient set of approximately independent extremes for stable inference across locations. The end-to-end pipeline from trajectory simulation to COR estimation is summarized in Fig. 2.5.
@@ 3 Extreme value theory (EVT)
Extreme Value Theory (EVT) provides the statistical foundation for estimating COR by extrapolating rare, high-severity events from frequent near-misses. In this study, extreme events are represented by block maxima of 2D-TTC derived from high-resolution vehicle trajectories (Algorithm 1). EVT enables risk estimation without requiring large volumes of historical crash data (e.g., Zheng and Sayed, 2020; Fu and Sayed, 2022a). Two EVT approaches are commonly applied: the Block Maxima (BM) method, which fits maxima or minima from fixed-duration blocks using the Generalized Extreme Value (GEV) distribution (Coles et al., 2001), and the Peaks Over Threshold (POT) method, which models threshold exceedances using the Generalized Pareto Distribution (GPD). Although POT can be data-efficient, it is highly sensitive to threshold selection and typically requires declustering and handling of dependencies when applied to dense trajectory data. By contrast, BM aligns naturally with real-time AV unique-segmented data, avoids subjective thresholds, and supports scalable time-series analysis (Fu and Sayed, 2022a; Songchitruksa and Tarko, 2006). Consequently, this study adopts the BM–GEV framework (see Fig. 2.5).
The GEV has been widely applied in traffic safety research. Early studies modeled near-misses using univariate and multivariate GEV formulations (e.g., citeporsini2020large, zhang2019real, zheng2014freeway). Recent advances incorporate nonstationary structures, linking GEV parameters to covariates such as vehicle dynamics (Anis et al., 2025b; Fu and Sayed, 2021c, 2022b). Bayesian hierarchical models have also addressed sparse data and unobserved heterogeneity across vehicles, sites, and traffic contexts. With increasing availability of AV trajectories and HD maps, there is growing motivation to combine geometry-aware near-miss extraction (including V–I exposure) with hierarchical EVT to support corridor-scale risk inference.
Despite this progress, most UGEV models still neglect the joint influence of fixed and time-variant parameters, which are critical but distinct sources of heterogeneity. For instance, (e.g., lane count, median type) is relatively stable within segments or intersections, whereas vehicle dynamics (e.g., speed, deceleration, spacing) vary within and across interaction windows and drive short-term volatility. Capturing this heterogeneity is crucial for estimating corridor or network-level crash risk. To address this gap, this study employs a hierarchical Bayesian structure grouped random parameters (HBSGRP) UGEV model. Unlike standard random-parameter models, which assign variation independently to each unit, grouped random parameters capture shared variation across clusters (e.g., segments, intersections) while allowing for within-group variability. This structure provides both flexibility and parsimony, making it well-suited for high-resolution, corridor-scale crash risk modeling (Washington et al., 2020; Cai et al., 2018; Islam et al., 2023).
@@ 3.1 Univariate generalized extreme value (UGEV)
The BM-based UGEV model is used to characterize the tail behavior of extreme (e.g., 2D-TTC), which represents the most critical events (near-miss: V–V and V–I interactions). To implement this framework, continuous vehicle trajectories are separated into fixed-duration windows. Within each window, the minimum 2D-TTC is extracted and negated to align with the upper-tail modeling convention (Coles et al., 2001). The resulting BM sequence serves as input for parameter estimation. Throughout this study, we focus on maxima, denoted by $M_{n}$, under the assumption that each block consists of $n$ independent and identically distributed (i.i.d., $X_{1},X_{2},\ldots,X_{n}$). To align with conventional UGEV formulations, we consider the negative values of the original variables derived from Algorithm 1, thereby defining the block maximum as $M_{n}=\max\{-X_{1},-X_{2},\ldots,-X_{n}\}$. Under suitable regularity conditions, if there exist sequences of constants $a_{n}\in\mathbb{R}$ and $b_{n}>0$, the normalized block maxima $M_{n}^{*}=(M_{n}-b_{n})/a_{n}$ converge in distribution to a non-degenerate limit $G(x)$ as $n\to\infty$, then: $\Pr\left(\frac{M_{n}-a_{n}}{b_{n}}\leq X\right)\xrightarrow[n\to\infty]{}f(x)$, where $G(x)$ is a non-degenerate distribution function. According to EVT, the limiting distribution $G(x)$ must belong to one of three families: Gumbel, Fréchet, or Weibull, each of which can be captured by the parametric form of the GEV distribution (Coles et al., 2001). Let the UGEV be defined as:
@@
3.1.1 Hierarchical Bayesian Structure Grouped Random Parameters (HBSGRP) model
To capture structured heterogeneity in extreme outcomes across the corridor, we employ the HBSGRP model. In this study, groups $k=1,\dots,K$ correspond to corridor elements (intersections or directional segments), so each block extreme $X_{k,i}$ is associated with a specific spatial context. Covariates are divided into (i) group-level fixed effects $Y$, which describe relatively stable roadway context (e.g., lane configuration, median type), and (ii) interaction-level dynamics $W$, which vary across blocks (e.g., relative speed, relative distance). The key feature of HBSGRP is that selected dynamic effects are allowed to vary by group, enabling systematic differences in interactions across corridor elements while maintaining corridor-wide coherence through partial pooling.
The model is formulated as a three-layer hierarchical Bayesian structure: a data layer that models block extremes using a UGEV distribution, a process layer that introduces grouped random parameters (partial pooling across corridor elements), and a prior layer that regularizes the model and quantifies uncertainty. Under a latent Gaussian formulation, the UGEV parameters are linked to covariates through identity link functions that combine global fixed effects with group-specific random effects, allowing the model to represent corridor-wide structure while preserving localized heterogeneity.
@@ 3.1.1 Hierarchical Bayesian Structure Grouped Random Parameters (HBSGRP) model
In this formulation, $q(\Theta\mid X)$ denotes the posterior distribution, integrating both the data likelihood and prior knowledge. The term $q_{\text{data}}(X\mid\Theta)$ corresponds to the likelihood function defined by the EVT assumptions. The component $q_{\text{process}}(\Theta\mid\Psi)$ captures the hierarchical dependency structure, including the group-varying parameters. Lastly, $q_{\text{prior}}(\Psi)$ encompasses the prior distributions of hyperparameters that control the model’s flexibility and penalize over-complexity. Eqs. (2329) therefore summarize the full HBSGRP–UGEV hierarchy used in this study.
@@
$q_{\text{data}}(X_{k,i}\mid\Theta)=\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}\frac{1}{\sigma_{k}}\exp\left\{-\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1-1/\xi_{k}}$ (23)
where $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$ are the location, scale, and shape parameters of the UGEV distribution for observation $i$ in group $k$. This layer expresses how block maxima of 2D-TTC are modeled as extreme outcomes that differ across groups and blocks.
Parameters are expressed as functions of group-level fixed-effect covariates $Y_{k,j}$ (e.g., lane no, lane width, median type, etc.) and sample-specific random-effect covariates $W_{k,i,j}$ (e.g., speed, acceleration, distance, etc.) as follows:
$\mu_{k,i}=\beta_{\mu,0}+\sum_{j=1}^{J}\beta_{\mu,j}Y_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\mu,j,k}W_{k,i,j}$ (24)
$\vartheta_{k,i}=\beta_{\sigma,0}+\sum_{j=1}^{J}\beta_{\sigma,j}Y_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\sigma,j,k}W_{k,i,j},\quad\sigma_{k,i}=\exp(\vartheta_{k,i})$ (25)
$\xi_{k,i}=\beta_{\xi,0}$ (26)
In these expressions, $Y_{k,j}$ denotes the $j$-th group-level fixed-effect covariate for group $k$, and $W_{k,i,j}$ denotes the $j$-th observation-level covariate for event $i$ within group $k$. The parameters $\beta_{\mu,0}$, $\beta_{\sigma,0}$, and $\beta_{\xi,0}$ are global intercepts for the location, log-scale, and shape parameters, respectively. The coefficients $\beta_{\mu,j}$, $\beta_{\sigma,j}$, and $\beta_{\xi,j}$ are global fixed-effect coefficients associated with geometric and static traffic variables, whereas $\gamma_{\mu,j,k}$ and $\gamma_{\sigma,j,k}$ are grouped random-effect coefficients that vary across groups $k$ and are associated with dynamic, time-varying covariates. Taken together, Eqs. (23)–(26) show how the HBSGRP model combines corridor-level geometric characteristics (through fixed effects on $Y_{k,j}$) with time-varying driving conditions (through group-specific random effects on $W_{k,i,j}$), making the formulation well suited for corridor-level risk analysis.
Therefore, the contribution of the process layer, modeling within-group variability, is defined as
@@
where $\bar{\mu}_{k}$, $\bar{\vartheta}_{k}$, and $\bar{\xi}_{k}$ are group-specific average parameters. The variance components $\tau_{\mu}^{2}$, $\tau_{\vartheta}^{2}$, and $\tau_{\xi}^{2}$ control the amount of between-group heterogeneity and induce shrinkage toward group means, thereby encoding the “grouped random parameters” aspect of the HBSGRP model and allowing different corridor elements to share information while retaining distinct risk profiles.
@@
where $\text{IG}(\alpha,\beta)$ denotes an inverse gamma prior distribution ensuring positivity constraints for variance parameters.These prior and hyperprior specifications provide regularization for both fixed and grouped random effects, thereby completing the hierarchical definition of the HBSGRP model.
@@ 3.2 COR estimation
Building on the HBSGRP–UGEV model, COR is quantified by mapping block-level extreme near-miss severity to the probability of a crash-level outcome. For each corridor element $k$ (intersection or directional segment), the trajectory stream is partitioned into fixed-duration interaction blocks, and the minimum 2D–TTC in each block is extracted and transformed to the UGEV scale (Section 3.1.1). These block extremes form a sequence $X_{k,i}$, which is modeled using a nonstationary UGEV distribution with parameters $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$. On the transformed scale, a crash-level event corresponds to the physical boundary at which the $\omega=0$ (i.e., 2D–TTC =0=0). Therefore, the probability of a crash occurring within the interaction window $i$ in group $k$ is defined as the upper-tail probability of the fitted UGEV distribution evaluated at $\omega=0$:
$P_{k,i}^{\text{crash}}=\Pr(X_{k,i}\geq\omega)=1-G(\omega;\mu_{k,i},\sigma_{k,i},\xi_{k,i})=1-\exp\left\{-\left[1+\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\},\omega=0$ (30)
Aggregating these block-level crash probabilities over an exposure period $T$ yields the expected COR for corridor element $k$:
$COR_{k}=\frac{1}{T}\sum_{i=1}^{N_{k}}P_{k,i}^{\text{crash}}$ (31)
where $N_{k}$ is the number of blocks observed for group $k$. The total expected crash frequency across the corridor is then obtained by summing over all groups:
$CF_{\text{total}}=\sum_{k=1}^{K}COR_{k}$ (32)
Negative TTC thresholds (e.g., 0.1-0.1 s to 3-3 s) are not interpreted as crash boundaries. Instead, they are used solely to probe increasingly severe regions of the near-miss distribution and to compare relative crash-prone risk across locations. Accordingly, COR should be interpreted as a model-based measure of crash-prone extreme conflict risk extrapolated from near-miss interactions, rather than as a direct count or prediction of observed crashes.